300 people were polled to see if a dog or cat were their favorite pet. 185 likes dogs 130 liked cats and 75 people liked both. how many people didnt like either pet?
What must i find?
What do i know?
What strategy should i use?


thank you very much for your help I greatly appreciate it

Answer by ankor@dixie-net.com(15623)

300 people were polled to see if a dog or cat were their favorite pet.
185 likes dogs 130 liked cats and 75 people liked both.
how many people didn't like either pet?
:
What must i find? Find out how many liked only dogs and how many liked only cats
185 - 75 = 110 liked dog only
130 - 75 = 55 liked cats only
:
What do i know?
Pet lovers consist of three groups, dogs only, cats only, & both
the total: 110 + 55 + 75 = 240
:
What strategy should i use?
Subtract pet lovers from 300
300 - 240 = 60 don't like either pet